A 500 N boy is standing on frictionless roller skates. If someone pulls him, horizontally, with 40 N of force to the North, what will be his acceleration?


1
Expert's answer
2021-01-20T18:00:48-0500

Let's first find the mass of the boy:


"m=\dfrac{W}{g}=\dfrac{500\ N}{9.8\ \dfrac{m}{s^2}}=51\ kg."

Then, we can find the acceleration of the boy from the Newton's Second Law of Motion:


"F=ma,""a=\dfrac{F}{m}=\dfrac{40\ N}{51\ kg}=0.78\ \dfrac{m}{s^2}."

Answer:

"a=0.78\ \dfrac{m}{s^2}."